What is y+3=4/5(x-7) in standard form?
Answer by Alan3354(69443)

What is y+3=4/5(x-7) in standard form?
Standard form is x and y terms on the left and no fractions
Multiply by 5
5y+15 = 4(x-7)
5y+15 = 4x-28
4x-28 = 5y+15
4x = 5y+43
4x - 5y = 43